PRESSURE BROADENING, SHIFTS, AND LINE MIXING IN THE ν_{4} BAND SYSTEM OF CH_{4}

\underline{MARY ANN H. SMITH}, Science Directorate, NASA Langley Research Center, Hampton, VA 23681-2199; D. CHRIS BENNER and V. MALATHY DEVI, Department of Physics, The College of William and Mary, Williamsburg, VA 23187-8795; A. PREDOI-CROSS, Department of Physics, The University of Lethbridge, Lethbridge, AB T1K 3M4, Canada.

A multispectrum nonlinear least squares technique^{a} has been used to examine self- and air-broadening, pressure-induced shifts, and line mixing in infrared spectra of methane (CH_{4}) in the 7-9 μm region. The temperature dependences of these parameters have also been determined for numerous transitions in the ν_{4} and ν_{2} bands of^{12}CH_{4}. The laboratory absorption spectra used in this study were recorded at high resolution (0.006-0.01 cm^{−1}) with the McMath-Pierce Fourier transform spectrometer of the National Solar Observatory. Sample temperatures ranged from 210 to 314 K, and broadening gas pressures were between 0.06 and 0.72 atm. The line broadening, shift and mixing parameters (off-diagonal relaxation matrix elements) were obtained by using the multispectrum technique to fit selected regions of 20 or more spectra simultaneously. In addition, accurate line center positions and absolute intensities were determined. Line mixing was observed in the Q branches and in the J-manifolds of the P and R branches of the ν_{4} bands of ^{12}CH_{4} and ^{13}CH_{4}. The mixing parameters resulting from the present study are compared to the results of other studies of line mixing in CH_{4}, and the effect of line mixing on the retrieved values of the line shifts will be discussed. This research is supported by NASA’s Upper Atmosphere Research Program.

aD. Chris Benner et al., JQSRT \underline{53}, 705-721 (1995).